(d) You buy a share of Ciao corporation stock for $21.40. You expect it to pay dividends of $1.07, and $1.1449 in Years 1, and 2 respectively. Required: (i) Calculate the growth rate in dividends. (2 marks)
media%2F839%2F839f916d-4281-4cb0-a858-c7
Answer #1

(i)

growth rate in dividends=1.1449/1.07-1=7.00%

(ii)

expected dividend yield=1.07/21.40=5.00%

(iii)

expected total rate of return=5.00%+7.00%=12.00%

the above is answer..